Please help to show the work. This is hard.
11. The total sales, S (in billions of dollars), for electronics and appliance stores in the United States for the years 1997 to 2004 are listed in the table below. Find a linear model that represents the data. (Hint: Please use the average of the slope of the first two points and the last two points). Let t represent the year, with t = 0 corresponding to 1997.
Year Sales
1997 64.9
1998 68.5
1999 70.2
2000 76.0
2001 81.9
2002 86.4
2003 85.2
2004 90.1

Answer by stanbon(75887)

Find a linear model that represents the data. (Hint: Please use the average of the slope of the first two points and the last two points). Let t represent the year, with t = 0 corresponding to 1997.
Year Sales
1997 64.9
1998 68.5
1999 70.2
2000 76.0
2001 81.9
2002 86.4
2003 85.2
2004 90.1
Points: (1997,64.9) and (2004,90.1)
slope = (90.1-64.9)/(7)= 25.2/7=18/5
#=mt+b
90.1=(18/5)(2004)+b
b=90.1-7214.4=-7124.3
EQUATION:
#=(3.6)t-7124.3
Cheers,
Stan H.